Detailed topic list for PartnerObserving Web

Results from PartnerObserving web retrieved at 21:20 (GMT)

12Apr2013_Log
0215 UT\xA0\xA0\xA0\xA0 Dome dark.\xA0 Mostly cloudy.\xA0 Beginning to take slitless pixel flats.0238 UT\xA0\xA0\xA0\xA0 Beginning MOS flatfield calibration for E...
12Apr2014_Log
April 11, 2014 19:30 Cloudy: Just done dark calibration and opening the telescope now. 19:38 Telluric acquisition HIP 035050, seeing ~1.4. Clouds 20:05 Lost 12...
12Jun2007
PartnerObserving Log for 2007 June 12/13 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue 02:38 Sunset 02:47 Dome Opened...
r7 - 21 Jul 2009 - 22:55 by NormCushing
12June2014_Log
June 12 Using MODS. * 19:50 opening, moving to bright star near M1115, clear, seeing is 1" * 20:14 start acquisition on M1115 * 20:29 start science on M1115 (2x20...
12Mar2007
PartnerObserving Log for 2007 Mar 12/13 Observers: Rick Pogge Inst. Support: John Hill TO: Ed Lundin UTC Mar 12 2100: Fired up the system, cryostat temperature ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
12Mar2011_Log
March 12, 2011 LUCIFER 1 Mask exchanger failed yesterday morning. Germans worked on it all day, but do not know what is wrong. We decided to do only imaging ton...
12May2007
PartnerObserving Log for 2007 May 12/13 MST Technical Observers: J. Hill, (A. DiPaola, R. Speziali) Telescope Operator: E. Lundin Instrument: LBC Blue Waiting for...
r4 - 21 Jul 2009 - 22:55 by NormCushing
12Oct2007
Observers: C.S. Kochanek, V. Testa, D. Gonzales HuertaClear sky but with a few thin clouds at horizon toward East. Opened at 01:00 01:10 Azimut encoders can't b...
13-14Feb2007
PartnerObserving Log for 2007 Feb 12/13 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Fog blowing over the summit...
r6 - 21 Jul 2009 - 22:55 by NormCushing
13Apr2007
$ 16:45: Biases and 900 sec Dark $ 18:30: The fog has lifted, but it's still mostly couldy, and humidity is 98%. The satellite images look promising, though. ...
r6 - 14 Apr 2007 - 11:02 by JasonHarris
13Apr2013_Log
0037 UT\xA0\xA0\xA0 Wake MODS take test frame0127 UT\xA0\xA0\xA0 Take bias frames0203 UT\xA0\xA0\xA0 Begin opening\xA0 Moderately high winds (10 15 m/s)0225 UT\...
13Apr2014_Log
April 12, 2014 19:09 Pointing to telluric, clouds are clearing, winds are high (15 m/s) 19:38 Telluric acquisition HIP35050, avg seeing 1.9 19:52 Telluric acquisi...
13June2007
PartnerObserving Log for 2007 June 13/14 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue 23:50 initialized camera! same...
r6 - 21 Jul 2009 - 22:55 by NormCushing
13Mar2007
PartnerObserving Log for 2007 Mar 13/14 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 13 2100: LBC shutter is still unrepo...
r6 - 21 Jul 2009 - 22:55 by NormCushing
13Mar2011_Log
March 13, 2011 Observers: J. Bechtold, D. Thompson, F. Bian, L. Jiang Cloudy all afternoon 6:50pm. Sun has set. Completely overcast. 7:05pm Started two 15 min...
13Oct2007
Observers: C.S. Kochanek, V. Testa, D. Gonzales Huerta Clouds all around. Humidity 80% at sunset (00:56) 00:54 5bias 005405 005625 01:00 Slew to 1hr E, 2...
14-15Feb2007
PartnerObserving Log for 2007 Feb 13/14 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Serious winter storm all af...
r7 - 21 Jul 2009 - 22:55 by NormCushing
14Apr2007
$ 16:15: 25 Bias frames and 900 sec Dark $ 18:45: Sky is mostly clear. Some wispy clouds in the northwest, so it's not photometric. $ 19:00: Doing Twilig...
r6 - 15 Apr 2007 - 12:20 by JasonHarris
14Apr2013_Log
UT date May 14 2013. Observer: F Bian, J Teske. Support Astronomer: M. Edwards Telescope Operator: Steve A Condition: some clouds around the horizon. No photometr...
14Apr2014_Log
April 13 10:50 Opening the dome after waiting for a high wind to subside. 11:03 Starting with A851. 11:20 Having trouble identifying the field. Something's wrong ...
14Feb2007
PartnerObserving Log for 2007 Feb 14/15 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Winter storm continues...Ac...
r6 - 21 Jul 2009 - 22:55 by NormCushing
14Jun2007
PartnerObserving Log for 2007 June 14/15 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: Very thin ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
14Mar2007
PartnerObserving Log for 2007 Mar 14/15 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 14 1745: Clock board from the CCD co...
r9 - 21 Jul 2009 - 22:55 by NormCushing
14Mar2011_Log
March 14, 2011 Observers: J. Bechtold, D. Thompson, F. Bian, L. Jiang 2:25pm. Lights off in the dome. Started series of 900s darks for Linhua's observations ton...
14Oct2007
Main.JohnHill 15 Oct 2007 Hi Matt Coleman and Dave Sand are at the LBT. Olga is helping and it is great. 0110 UT Opening Dome. Having some trouble retracting....
15Apr2007
$ 18:50: Welcome to LBT, Dr. Sand! High winds, but it should be okay. $ 19:10: Twilight flats Filter scale counts U 1.0 30000, 25000, * V ...
r3 - 16 Apr 2007 - 09:37 by JasonHarris
15Apr2013_Log
Observer: Fuyan Bian, Cai Zheng, Ran Wang. Condition: Cloudy in the first half night. 08:37 we open the telescope, go to zheng's targets mask1_PA50, something wro...
15Apr2014_Log
April 14 7:00 Clear sky. 7:35 Moving to telluric star HIP53735. Decent seeing 0.8 0.9". 7:45 Start taking H and K band telluric star spectra. 7:55 Moving to CLJ1...
15Apr2015_Log
Apr 15 (UT) Observer: Wen fai Fong 02:25 UT 8 deg twilight 1.5 hr of time lost due to clouds 03:55 UT Open telescope, 0.8 1.2" seeing 04:22 UT telescope clo...
15Feb2007
PartnerObserving Log for 2007 Feb 15/16 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill, Dave Thompson TOs: Ed Lundin David Gonzalez Beautifully ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
15Jun2007
PartnerObserving Log for 2007 June 15/16 Observers: V. Burwitz, J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: Thin high clou...
r8 - 21 Jul 2009 - 22:55 by NormCushing
15Mar2007
PartnerObserving Log for 2007 Mar 15/16 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 15 2200: Day crew is replacing the s...
r10 - 21 Jul 2009 - 22:55 by NormCushing
15Oct2007
PartnerObserving Log for 2007 Oct 15 16 Observers: D. Sand and M. Coleman Telescope Operator: D. Gonzalez Huerta Beautiful Clear evening. Not a cloud in the sky....
r7 - 21 Jul 2009 - 22:55 by NormCushing
15Sep2007
PartnerObserving Log for 2007 Sep 15 16 Observers: J. Hill, O. Kuhn Telescope Operator: A. Ceranski LBTO Family night, but cloudy until midnight. Ray and Tim and ...
r4 - 21 Jul 2009 - 22:55 by NormCushing
16Apr2007
PartnerObserving Log for 2007 April 16/17 Observers: J. Harris, D. Sand Instrument Support: Telescope Operator: Instrument: LBC Blue $ 19:00: Camera shutter ...
r5 - 21 Jul 2009 - 22:55 by NormCushing
16Apr2012_Log
Observers: Linhua Jiang, Emily McLinden 7:20 pm: technical tests by Olga. 8:00 pm: pointing to Emily's target obj1. 8:20 pm: struggling with guide star, turned ou...
16Apr2013_Log
Time: MST 8:00pm 9:15pm Stars are very elongated even in the center of the field. Science operation team strongly suspects this is the secondary mirror problem...
16Apr2014_Log
Apr 15 6:20 Starting up LUCI on the remote control room at Steward. 7:26 Start collimating the telescope. 7:35 Collimation finished. Seeing=0.85" 7:35 Going to te...
16Apr2015_Log
Apr 16 (UT) Observer: Wen fai Fong Did not open tonight due to high winds and dust. TO and observer collectively called the night 10:45 UT (40 minutes until twili...
16Feb2007
PartnerObserving Log for 2007 Feb 16/17 Observer: Tom Herbst, Giacomo Beccari Inst. Support: Dave Thompson TOs: David Gonzalez, Aaron Ceranski 18:05 Mostly cle...
r8 - 21 Jul 2009 - 22:55 by NormCushing
16Jun2007
PartnerObserving Log for 2007 June 16/17 Observers: V. Burwitz, J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: thin clouds ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
16Mar2007
PartnerObserving Log for 2007 Mar 16/17 Observers: Rick Pogge, Peter Garnavich Support: John Hill on site, Olga Kuhn on call TO: Aaron Ceranski Afternoon weather...
r9 - 21 Jul 2009 - 22:55 by NormCushing
16Oct2007
PartnerObserving Log for 2007 Oct 16 17 Observers: D. Sand and M. Coleman Telescope Operator: A. Ceranski Main.JohnHill 17 Oct 2007 Sand and Coleman here. ...
r5 - 21 Jul 2009 - 22:55 by NormCushing
16Sep2007
19:45 Stop and restart firefox on lbtmu03, since it was taking up to 62% memory and was slow to respond. In new firefox, start LBC and run up. 19:50 Start series ...
17Apr2007
PartnerObserving Log for 2007 April 17/18 Observers: J. Harris, D. Sand Instrument Support: D. Thompson, R. Bertram Telescope Operator: Instrument: LBC Blue $...
r6 - 21 Jul 2009 - 22:55 by NormCushing
17Apr2012_Log
Observers: Linhua Jiang, Emily McLinden 6:40 pm: start luci. 7:25 pm: slew to Emily's target obj1, and do acquisition. 7:47 pm: start observing obj1 (clear sky an...
17Apr2014_Log
Apr 16 7:30 starting collimation, wind West ~10 15m/s, clear 7:35 starting telluric HIP35050 for MOS in J band with 210_zJHK. Waiting for collimation. Facing the ...
17Feb2007
Observer1, Nicolas Bouche, Peter Buschkamp Observer2, Giacomo Beccari Inst. Support, Dave Thompson TOs; David Gonzalez, Aaron Ceranski Nice sunset (means lots of ...
17Jan2013_Log
Main.Walth 18 Jan 2013 We started out clear at 1" seeing, which improved to sub arcsec for most of the night. Near the end of the A611 observation we noticed thi...
Number of topics: 50
< Previous Page 2 of 32 Next >

See also the faster WebTopicList
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback